John had $30,000. He invested part of his money in bonds paying 12% annual simple interest and the rest into a savings account giving 4% annual interest. At the end of the year he received $2,400 as extra income. How much did John place in each investment?
Answer by mananth(16946)

Bonds 12.00% per annum ------------- Amount invested =x
Savings 4.00% per annum ------------ Amount invested = y
20000
Interest----- 740.00

Bonds 12.00% per annum ---x
Savings 4.00% per annum ---y
Total investment
x + 1 y= 20000 -------------1
Interest on both investments
12.00% x + 4.00% y= 740
Multiply by 100
12 x + 4 y= 74000.00 --------2
Multiply (1) by -12
we get
-12 x -12 y= -240000.00
Add this to (2)
0 x -8 y= -166000
divide by -8
y = 20750
Bonds 12.00% $ -750
Savings 4.00% $ 20750

CHECK
-750 --------- 12.00% ------- -90.00
20750 ------------- 4.00% ------- 830.00
Total -------------------- 740.00

m.ananth@hotmail.ca